OpenStudy (anonymous):

Parker earned $32 by selling 27 glasses of fresh-squeezed lemonade at his lemonade stand on a Saturday morning. He charged $1.00 for a small glass and $1.50 for a large glass. Use a model to determine how many large glasses he sold.

OpenStudy (anonymous):

Let large glass be x, small glass be y then x + y = 27 1.5x + y = 32 solve for x and y

OpenStudy (anonymous):

0.5x = 5 x = 5/0.5 x = 10 y = 17

OpenStudy (anonymous):

oh! thank you! i get it!!